What type of good is this commodity? 9. A researcher estimated that the price elasticity of demand for automobiles in the United States is -1.2, while the income elasticity of demand is 3.0. Next year, U.S. automakers intend to increase the average price of automobiles by 5 percent, and they expect consumers disposable income to rise by 3 percent. (a) If sales of domestically produced automobiles are 8 million this year, how many automobiles do you expect U.S. automakers to sell next year? (b) By how much should domestic automakers increase the price of automobiles if they wish to increase sales by 5 percent next year?
Answer #1

Price elasticity of demand=ep=-1.2

Income elasticity of demand=ei=3.0

Change in average price=+5%

Change in disposable income=+3%

Current level of sales of automobiles=q=8 million

a)

Expected sale next year=q* (1+Change in average price*ep)*(1+Change in disposable income*ei)

=8*(1+5%*(-1.2))*(1+3%*3.0)

=8.1968

b)

Expected change in sales=5%

Expected change in average price=x

Change in disposable income=3%

Change in sales=(1+Change in average price*ep)*(1+Change in disposable income*ei)-1

5%=(1+x*(-1.2))*(1+3%*3)-1

1.05=(1-1.2x)*1.09

0.963303=1-1.2x

0.963303=1-1.2x

x=0.036697 or say 3.67%